OpenStudy (anonymous):

simplify radical 8a^4 b^9

OpenStudy (agreene):

\[\sqrt{8a^4b^9}\] You can separate the terms to: \[\sqrt{8a^4}\sqrt{b^9}\] then simplify each: \[\sqrt[3]{2} a^2 \sqrt{b^9}\] or 2^(3/2) a^2 sqrt(b^9) if you prefer.